OpenStudy (anonymous):

I drive to a house at 40 mph and drive back at 45 mph, the trip took 4 hours how long was my trip?

OpenStudy (anonymous):

Distance=speed*time

OpenStudy (bahrom7893):

wait what about the trip in one direction?

OpenStudy (anonymous):

40x=45y x+y=4 x = time taken to go the house y = time taken to return

OpenStudy (anonymous):

x=4-y substitute this in 40x=45y